OpenStudy (anonymous):

Solve the differential equation y''+2y'+2y=sin(t)e^(-t) I'm mainly having problems with the practical solution and wondering if there is a better way of doing it. The practical solution i'm thinking is y(prac)=t*e^(-t)*(A*cos(t)+B*sin(t)) Will i need to take derivatives and put back into the original equation then solve or is there a better method for doing this ?

OpenStudy (anonymous):

I think \(y_p=e^{-t}(A\cos t+B\sin t)\) would work.

OpenStudy (anonymous):

solve your characteristic equation, good so your solution will be \[y=-1/2 e^{-t}tcos(t) c1e^{-t}\cos(t)+c2e^{-t}\sin(t) \]

OpenStudy (anonymous):

m^2+2m+2 = -1+/-i

OpenStudy (anonymous):

Or not. It's not right !

OpenStudy (anonymous):

@Anwara it has to have t as well as it is the same as the complimentary equation otherwise. I'm stuck solving the (Acos(t) + Bsin(t)) in the y(prac). I'm wondering if i need to take derivatives of the y(prac) and then put this back into my original equation and solve for A and B or if there is a better way. As the derivatives are rather long

OpenStudy (anonymous):

@JohnnyBreur sorry i should have been more clear. The issue i'm having is solving the A and B in the y(p) must i take the derivatives and put them back into the original equation and solve for A and B or is there a better method

OpenStudy (anonymous):

back sub to solve for A and B

OpenStudy (anonymous):

you start with y''+2y'+2y=0n find your solution

OpenStudy (anonymous):

Therefore i have to solve for y'', y' of the prac to back sub and find my solution. e.g. \[y_p'=\exp(-t)(-A t \sin(t)+A \cos(t)-A t \cos(t)+B \sin(t)-B t \sin(t)+B t \cos(t))\] And solve y'' and then insert into the equation? I was just trying to find a method that wouldnt involve such a large equation if there were one otherwise i'll just have to do it this way

OpenStudy (anonymous):

Do you know the annihilator approach? It might be easier to use it in this case.

OpenStudy (anonymous):

No i'm unfamilar with the annihilator approach did you have a good link to info on it? Might be worth a look

OpenStudy (anonymous):

i think you need an annihilator here (D^2-2aD+a^2+b^2) this will annihilate e^atsin(bt) since a=(-1) b=1 (D^2-2(-1)D+1+1) sin(t)e^(-t) D=cos(t)e^-t-sin(t)e^-t D^2=-sin(t)e^-t-cos(t)e^-t-cos(t)e^-t+sin(t)e^-t 2D=(2cos(t)e^-t-2sin(t)e^-t) D^2=(-2cos(t)e^-t THE annihilator does not look like its going to 0 (-2cos(t)e^-t+2) you think i went wrong somewhere ANWARA?